OpenStudy (anonymous):

simplify (t^2+6t/t-1)-(7/t-1)

OpenStudy (asnaseer):

both fractions have the same denominator, so you should be able to subtract the numerators directly.

OpenStudy (asnaseer):

e.g.:\[\frac{a}{b}-\frac{c}{b}=\frac{a-c}{b}\]

OpenStudy (anonymous):

so then i should come up with T^2+6t+7/t-1? right ?

OpenStudy (asnaseer):

not quite - you have added the fractions instead of subtracting them

OpenStudy (anonymous):

so then it would -7 instead.

OpenStudy (asnaseer):

yes - and now you should be able to factorise the numerator in:\[\frac{t^2+6t-7}{t-1}\]

OpenStudy (anonymous):

i have to do grouping in order to factor , but what do i do with the denominator ?

OpenStudy (asnaseer):

just concentrate on the numerator for now - see if you factorise: \(t^2+6t-7\)

OpenStudy (anonymous):

i got (t+6)(t+1)?

OpenStudy (asnaseer):

that doesn't look quite right - that expands to: \(t^2+7t+6\)

OpenStudy (asnaseer):

you need to find two numbers that multiply to give -7 and add to give +6

OpenStudy (anonymous):

well idk what they would be. could it be -1 and + 7.

OpenStudy (asnaseer):

that is EXACTLY what they would be :)

OpenStudy (asnaseer):

so now you can use those to factorise the numerator as follows:\[t^2+6t-7=(t-1)(t+7)\]

OpenStudy (asnaseer):

understand so far?

OpenStudy (anonymous):

then i would just wind up with x+7 as an answer because the x-1 would canncel in the numerator and denominator.

OpenStudy (asnaseer):

perfect! but remember the variable here is 't' and not 'x' well done! :)

OpenStudy (anonymous):

ohhh duh ! im just used to using x . and thanks for the medal (:

OpenStudy (asnaseer):

you deserved it. :)

OpenStudy (anonymous):

can i bother you with another question?

OpenStudy (asnaseer):

sure - just post each question in the list on the left please

OpenStudy (anonymous):

okayy .